# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v1

@@ Abstract
The rapid advancements in computing dramatically increase the scale and cost of training Large Language Models (LLMs). Accurately predicting downstream task performance prior to model training is crucial for efficient resource allocation, yet remains challenging due to two primary constraints: (1) the “emergence phenomenon”, wherein downstream performance metrics become meaningful only after extensive training, which limits the ability to use smaller models for prediction; (2) Uneven task difficulty distributions and the absence of consistent scaling laws, resulting in substantial metric variability. Existing performance prediction methods suffer from limited accuracy and reliability, thereby impeding the assessment of potential LLM capabilities. To address these challenges, we propose a Clustering-On-Difficulty (COD) downstream performance prediction framework. COD first constructs a predictable support subset by clustering tasks based on difficulty features, strategically excluding non-emergent and non-scalable clusters. The scores on the selected subset serve as effective intermediate predictors of downstream performance on the full evaluation set. With theoretical support, we derive a mapping function that transforms performance metrics from the predictable subset to the full evaluation set, thereby ensuring accurate extrapolation of LLM downstream performance. The proposed method has been applied to predict performance scaling for a 70B LLM, providing actionable insights for training resource allocation and assisting in monitoring the training process. Notably, COD achieves remarkable predictive accuracy on the 70B LLM by leveraging an ensemble of small models, demonstrating an absolute mean deviation of 1.36% across eight important LLM evaluation benchmarks.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as transformative technologies in natural language understanding, generation, and reasoning [1, 14, 5]. Their impressive success heavily relies on scaling model parameters and pre-training data, with training loss empirically following a power-law relationship with compute [18, 21]. However, this reduction in training loss primarily reflects an in-domain compression effect and does not necessarily indicate improved out-of-domain generalization or downstream performance–the factor of primary concern in practice. Specifically, performance scaling of downstream tasks aims to predict the accuracy of the target LLM on downstream tasks using metrics from smaller models. Our objective is to develop a prediction method that works reliably across a diverse range of downstream tasks, minimizing the worst-case prediction error.
@@
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate large-model performance by modeling the performance-loss relationship [6, 13, 8, 38, 26], but this often fails to capture the emergent behaviors of LLMs and the mismatch between in-domain loss and downstream metrics [42]. Another line of research focuses on direct extrapolation of performance-compute relationship [1, 19], yet the uneven difficulty distribution across different evaluation samples undermines its accuracy. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give the detailed analysis in Section 3.
@@
To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [21], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns. Building on the performance scaling law, we develop a Clustering-On-Difficulty (COD) multi-stage framework for predicting downstream performance. Specifically, we first cluster tasks by their difficulty features, and then filter out clusters that lack valid extrapolation patterns. Next, we fit the performance-compute relationships in the remaining clusters under our performance scaling law, extrapolate the performance of large models within these clusters, and finally map the aggregated predictions to the complete task set.
@@
We validate our COD approach on eight evaluation sets, including popular MATH [15], BBH [32], and MMLU pro [36]. COD achieves an average prediction error of 1.36% on a 70B-parameter LLM. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
Our contributions can be summarized as follows:
@@
  • We propose the COD framework to address high variance and emergent phenomena in the LLM performance by effectively modeling the difficulty distribution within the evaluation sets.
@@
  • Extensive experiments conducted across eight diverse evaluation sets demonstrate that COD achieves a state-of-the-art average prediction error of 1.36% on a 70B-parameter LLM.
@@ 2.1 Loss Scaling Laws
Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the final performance of LLMs. Early work by Kaplan et al. [21] demonstrates that the pre-training loss of LLMs follows a power-law relationship with the compute (the number of floating-point operations) used in training. Subsequent studies extend these findings to other domains, such as computer vision [41], graph learning [24] and vision-language models [2, 16]. Recent research has also explored scaling laws in specific contexts, such as fine-tuning [17, 34], vocabulary size optimization [33], retrieval-augmented models [30], and hyperparameter tuning [23, 40]. These studies highlight the broad applicability of scaling laws and their potential to guide the efficient allocation of computational resources.
@@ 2.2 Downstream Task Performance Scaling
1. Loss-intermediate prediction. These methods predict the final training loss (or in-domain validation loss) of LLMs with loss scaling laws first, and then predict downstream performance through loss-performance relationships [6, 13, 8]. While these methods leverage established scaling laws for loss predictions, they encounter a fundamental limitation: the inconsistent mapping between loss and performance metrics. In addition, Xiao et al. [38] employ the evaluation set answer loss as an intermediate variable for estimation. Although answer loss correlates with the final performance metrics, its predictability remains low as predicting answer loss shares the challenges with predicting performance, including emergence phenomenon and high variance in task difficulty.
2. End-to-end performance-compute prediction. These methods [19, 26, 1] directly model the relationship between performance and compute (or the number of model parameters). Additionally, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set. Hu et al. [19] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
@@ 3 Pilot Study
Figure 1: Performance-loss relationship across difference model size (left) and learning rate schedule (middle). At equivalent loss values, smaller models or those with lower learning rates generally achieve higher accuracy than larger models or those with higher learning rates. Performance-compute relationship for different clusters of BBH samples(right). Different clusters demonstrate diverse scaling patterns.
@@
Training loss may mismatch downstream tasks performance. Predicting downstream performance based on training loss relies on the assumption that LLMs achieve identical downstream performance at the same loss value–an assumption that often does not hold. In practice, training loss primarily serves as an indicator of in-domain fitting, whereas downstream tasks typically represent out-of-domain evaluations. Moreover, training configurations, such as model size and learning rate, can significantly affect not only the final loss but also the model’s generalization capabilities.
@@
Diverse scaling patterns within the evaluation set. Scaling patterns capture the performance–compute relationship for a single task sample. However, different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or set of fitting functions) that generalizes well across diverse task samples. Fig. 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [32], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream benchmark.
@@
Figure 2: The pipeline of Cluster-On-Difficulty downstream task performance scaling, including 4 stages: a. Represent task difficulty feature with task-wise passrate vector. Cluster on difficulty feature and filter outliers. b. Fit cluster-wise performance-compute curve. Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent cluster. c. Predict accuracy on extrapolatable clusters. d. Map subset accuracy prediction to full evaluation set performance.
@@ 4 Method
In this section, we introduce the COD method in four parts, illustrated in Fig. 2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Section 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which enhances the benefit of extrapolating the performance-compute relationship for task clusters with similar difficulty features (Section 4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Section 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Section 4.4).
@@ 4.1 Clustering on Difficulty
Despite sharing common themes, tasks within evaluation sets demonstrate substantial difficulty differences. These differences result in diverse performance scaling patterns across tasks, making it challenging to apply a universal fitting function for predictions. Instead, we propose clustering tasks with comparable performance scaling behaviors to enable more accurate predictions. This approach minimizes the heterogeneity of difficulty features within clusters while ensuring that each cluster contains a sufficient number of samples for robust evaluation.
We adopt the passrate metric to quantify the capabilities of small-scale models [19]. For each model, we conduct multiple evaluation runs (e.g., 100 trials) and calculate the mean accuracy as the expected probability of correct responses. For each task, we characterize its difficulty through the passrates of models of increasing size. These passrates are arranged in ascending order of model scale, forming feature vectors that ideally exhibit monotonic growth within the [0, 1] range, as model capability typically increases with size. However, we observe that some tasks deviate from the expected scaling pattern, showing non-monotonic difficulty features. This phenomenon may be attributed to metric instability or fluctuations in model performance during training.
@@
Among classical clustering algorithms, the K-Means algorithm [25] needs to specify the number of clusters in advance. Although there are methods for automatically selecting the optimal number of clusters, e.g. Elbow method [35] and Silhouette [27], these methods need to introduce additional hyperparameters. DBSCAN [11] is a non-parametric density-based clustering algorithm that marks points in low-density regions as outliers while cluster points in the connected high-density regions. In practice, DBSCAN may lead to a larger final intra-class variance and does not meet the clustering requirements of the current task. MeanShift [12] algorithm adopts an extra clustering radius parameter to constrain the intra-class variance, which better fits our demands.
Figure 3: t-SNE visualization of different clustering methods. Each point represents an evaluation sample. DBSCAN(left): Continuous diffusion of clustering leads to too many samples within a class and large inner-group variance. MeanShift(Middle): Keep the high-density region as a unique group. Improved-MeanShift(Right): Constrain inner-group variance with the radius parameter.
To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of evaluation tasks on BBH [32]. Each point represents an evaluation sample and its color denotes the cluster type. Fig. 3. The improved MeanShift prevails, as it effectively splits dense areas into reasonable clusters. We explain the details of clustering algorithms in Section A.1, and smoothing techniques in Section A.2.
@@ 4.2 Fitting
Following cluster analysis, we compute evaluation metrics of small models within each cluster and conduct separate extrapolation curve fitting procedures. Small models are trained with the same ratio of training tokens to Compute Per Token (CPT). We propose a scaling law for downstream task performance, supported by theoretical analysis, which allows us to derive prediction formulas for performance scaling within clusters of tasks that share similar difficulty features. The fitting process initially excludes outlier samples, focusing only on the clustered sample set. For each cluster identified in the previous step, we compute accuracy metrics across small models, yielding an expected accuracy array for each cluster. By fitting these accuracy values against the computational costs of small models, we derive the expected accuracy-to-compute curve for each cluster.
We derive the fitting formula for the downstream task scaling law based on the following three assumptions:
@@
    The relationship between the answer loss and the compute follows a power law, which generalizes the power law in loss prediction into (Question, Answer) format data.
@@
    For task samples with a finite set of answers, the model gives a random guess choice if it cannot accurately solve it.
@@
    The task passrate is defined as the product of the predicted probabilities for each token, implying that each task sample has a unique answer, and the model outputs the answer only without any intermediate reasoning progress.
Note that these assumptions may not perfectly hold in practice, we provide additional discussions on Assumption 3 in Section 6. Under the above assumptions we can derive the scaling law for downstream task performance.
Proposition 1 (Scaling Law for Downstream Task Performance).
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions: The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu)$
where:
  • $g$ represents the random guess performance floor;
  • $1-g$ represents the maximum achievable performance improvement;
  • $a,b,c$ are positive constants;
  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$;
  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
We outline the key proof intuition here, with detailed proofs provided in Appendix B. Like existing approaches [19], we aim to establish the relationship between $\mathrm{loss}_{\mathrm{ans}}$ and model $passrate$, leveraging loss power-law scaling to derive a scaling formula for downstream task passrate metrics.
Proof intuition
The assumption 3 ensures unique task answers and neglecting the impact of model thinking before answers, the probability of correct task completion equals the product of token probabilities in the model output. This implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and $passrate$ for individual tasks.
Previous works overlook that computing the $passrate$ metric for an evaluation set requires averaging $\exp(-\mathrm{loss_{\mathrm{ans}}})$ across tasks, whereas applying the loss scaling law necessitates averaging loss before exponentiation. Mathematically, the performance scaling law computes the arithmetic mean of $\exp(-\mathrm{loss_{\mathrm{ans}}})$, while the loss scaling law after exponentiation yields the geometric mean.
We show that the difference between arithmetic and geometric means can be estimated by $\sigma^{2}/2\mu$, where $\sigma^{2}$ and $\mu$ denote the variance and mean of task passrates, respectively. Consequently, the downstream tasks performance scaling law derived from the loss scaling law is valid only for evaluation sets with limited difficulty variance. Our proposed clustering-based COD method constrains the variance of difficulty features within clusters, enabling better alignment with the performance scaling law.
Finally, we constrain the model output space to a finite answer set, random guessing yields an expected score $g$ for unsuccessful attempts.
Proposition 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
@@
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for the model on this cluster. Parameters $a$, $b$, $c$, and $g$ are to be fitted.

### v2

@@ Abstract
The escalating scale and cost of Large Language Models (LLMs) training necessitate accurate pre-training prediction of downstream task performance for efficient resource allocation. This is challenged by: 1) the emergence phenomenon, where metrics become meaningful only after extensive training, hindering prediction by smaller models; and 2) uneven task difficulty and inconsistent performance scaling patterns, leading to high metric variability. Current prediction methods lack accuracy and reliability. We propose a Clustering-On-Difficulty (COD) framework for downstream performance prediction. The COD framework clusters tasks by their difficulty scaling features, thereby establishing a more stable and predictable support subset through the exclusion of tasks exhibiting non-emergent behavior or irregular scaling. We adopt a performance scaling law to predict cluster-wise performance with theoretical support. Predictable subset performance acts as an intermediate predictor for the full evaluation set. We further derive a mapping function to accurately extrapolate the performance of the subset to the full set. Applied to an LLM with 70B parameters, COD achieved a 1.36% average prediction error across eight key LLM benchmarks, offering actionable insights for resource allocation and training monitoring of LLMs pretraining.
[…]
Chenyin Xu at , Ke Shen at
@@ 1 Introduction
Large Language Models (LLMs) have emerged as transformative technologies in natural language understanding, generation, and reasoning [1, 15, 5]. Their impressive success heavily relies on scaling model parameters and pre-training data, with training loss empirically following a power-law relationship with compute [19, 22]. However, this reduction in training loss primarily reflects an in-domain compression effect and does not necessarily indicate improved out-of-domain generalization or downstream performance–the factor of primary concern in practice. Specifically, performance scaling of downstream tasks aims to predict the accuracy of the target LLM on downstream tasks using metrics from smaller models. Our objective is to develop a prediction method that works reliably on a diverse range of downstream tasks, optimizing the worst-case prediction error.
@@
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [6, 13, 8, 38, 28], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [43]. Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 20], yet the uneven difficulty distribution between different evaluation samples undermines its accuracy.
A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give a detailed analysis in Sec. 3.
@@
To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [22], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns. Building on the performance scaling law, we develop a Clustering-On-Difficulty (COD) multi-stage framework for predicting downstream performance. Specifically, we first filter out clusters that lack scaling properties using an improved MeanShift clustering algorithm. Next, we fit the performance-compute relationships in the remaining clusters under our performance scaling law, extrapolate the performance of large models within these clusters, and finally map the aggregated predictions to the complete task set.
@@
We validate our COD approach on eight popular evaluation sets, including MATH [16], BBH [33], and MMLU pro [36] datasets. COD achieves an average prediction error of 1.36% on an LLM with 70B parameters. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
Our contributions can be summarized as follows.
@@
  • We propose the COD framework to address high variance and emergent phenomena in LLM performance scaling by effectively modeling the difficulty distribution within the evaluation sets.
@@
  • Extensive experiments conducted in eight different evaluation sets demonstrate that COD provides reliable predictions with an average prediction error of 1. 36% on an LLM with 70B parameters.
@@ 2.1 Loss Scaling Laws
Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the LLM performance. Early work by Kaplan et al. [22] demonstrates that the pre-training loss of LLMs follows a power-law relationship with the compute (the number of floating-point operations) used in training. Subsequent studies extend these findings to other domains, such as computer vision [42], graph learning [26], and vision-language models [2, 17]. Recent research has also explored scaling laws in specific contexts, such as fine-tuning [18, 35], vocabulary size optimization [34], retrieval-augmented models [31], and hyperparameter tuning [24, 41]. These studies highlight the broad applicability of scaling laws and their potential to guide the efficient allocation of computational resources.
@@ 2.2 Downstream Task Performance Scaling
Loss-intermediate prediction. These methods predict the final training loss (or in-domain validation loss) of LLMs with loss scaling laws first, and then predict downstream performance through loss-performance relationships [6, 13, 8]. While these methods leverage established scaling laws for loss predictions, they encounter a fundamental limitation: the inconsistent mapping between loss and performance metrics. In addition, Xiao et al. [38] employ the evaluation set answer loss as an intermediate variable for estimation. Although answer loss correlates with the final performance metrics, its predictability remains low as predicting answer loss shares the challenges with predicting performance, including emergence phenomenon and high variance in task difficulty.
End-to-end performance-compute prediction. These methods [20, 28, 1] directly model the relationship between performance and compute (or the number of model parameters). Additionally, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set. Hu et al. [20] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
@@ 3 Pilot Study
Figure 1: Performance-loss relationship across different model sizes (left) and learning rate schedules (middle). Performance-compute relationship for different clusters of the BBH samples(right)
@@
Training loss may mismatch downstream task performance. Predicting downstream performance from training loss assumes LLMs achieve identical downstream results at the same loss value, which does not hold universally. In practice, training loss primarily serves as an indicator of in-domain fitting, whereas downstream tasks typically represent out-of-domain evaluations. Moreover, training configurations, such as model size and learning rate, can significantly affect not only the final loss but also the model’s generalization capabilities.
@@
Diverse scaling patterns within the evaluation set. Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples. Fig. 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [33], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
@@
Figure 2: The pipeline of Cluster-On-Difficulty downstream task performance scaling, including 4 stages: a. Represent task difficulty feature with task-wise passrate vector. Cluster on the difficulty feature and filter outliers. b. Fit cluster-wise performance-compute curve. Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent clusters. c. Predict accuracy on extrapolatable clusters. d. Map subset accuracy prediction to full evaluation set performance.
@@ 4 Method
In this section, we introduce the COD method in four parts, as illustrated in Fig. 2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Sec. 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec. 4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Sec. 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec. 4.4).
@@ 4.1 Clustering on Difficulty
Tasks within evaluation sets, despite common themes, show substantial difficulty differences, leading to diverse scaling patterns that hinder universal fitting functions. We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster. To quantify task difficulty, we leverage the passrate(mean accuracy across multiple trials) from small models of progressively increasing scales. This process generates a feature vector for each task, which ideally exhibits monotonic growth, thereby reflecting the incremental enhancement of model capabilities.
@@
While K-Means [27] requires pre-specifying cluster numbers and DBSCAN [11] may yield large intra-cluster variance, MeanShift [12] allows constraining this variance via a radius parameter, better suiting our needs.
Figure 3: t-SNE visualization of different clustering methods: DBSCAN(left), MeanShift(Middle), Improved-MeanShift(Right). Each point represents an evaluation sample.
To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of evaluation tasks on BBH [33]. Each point represents an evaluation sample, and its color denotes the cluster type. As shown in Fig. 3, our improved MeanShift effectively splits dense areas. We explain the details of clustering algorithms in Sec. A.1, and smoothing techniques in Sec. A.2.
@@ 4.2 Fitting
After clustering, we compute metrics for small models within each cluster. These small models are trained with the same ratio of training tokens to compute per token. We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers. We derive the following fitting formula for the downstream task scaling law based on the training loss scaling law.
Theorem 1 (Scaling Law for Downstream Task Performance).
@@
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
@@
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
@@
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (1)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
Proof Sketch.
Assumption 3 (unique answers, no pre-answer reasoning) implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and individual task passrate, as correct task completion probability becomes the product of answer token probabilities. A key distinction, often overlooked, is how averaging occurs: the passrate metric averages $\exp(-\mathrm{loss}_{\mathrm{ans}})$ (an arithmetic mean), whereas applying the loss scaling law effectively averages loss before exponentiation (yielding a geometric mean). We show that the difference between these means can be approximated by $\frac{\sigma^{2}}{\mu}$ (where $\sigma^{2}$ and $\mu$ are the variance and mean of task passrates, respectively). Consequently, a performance scaling law derived from the loss scaling law is most valid for evaluation sets with limited difficulty variance. Our COD method achieves this by clustering tasks, thereby constraining intra-cluster difficulty variance and enabling better alignment with the performance scaling law. Finally, for tasks with a finite answer set, a random guessing score $g$ accounts for expected performance on unmastered tasks. Detailed proofs are provided in Appendix B. ∎
theorem 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
@@
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster. Parameters $a$, $b$, $c$, and $g$ are to be fitted. Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix G.